A Space & astronomy forum. SpaceBanter.com

Go Back   Home » SpaceBanter.com forum » Space Science » Space Station
Site Map Home Authors List Search Today's Posts Mark Forums Read Web Partners

ISS On-Orbit Status, 24-04-2004

Thread Tools Display Modes
Old May 1st 04, 12:55 PM
Jacques van Oene
external usenet poster
Posts: n/a
Default ISS On-Orbit Status, 24-04-2004

ISS On-Orbit Status 24 Apr 2004

All ISS systems continue to function nominally, except those noted
previously or below. Day 4 of joint Exp.8/Exp.9 operations.

With wake-up at midnight (EDT) and sleeptime at 3:30pm, both crews had
another long day with a demanding schedule, after receiving kudos on their
outstanding job yesterday on four PAO events and the Lab Window procedure,
despite a timeline equally as strenuous as today's.

Handover activities between Exp. 8 & Exp. 9 continued to go well. There
were 30 min. allotted on today's schedule for general handover, performed by
Michael Fincke and the Expedition 8 crew, in addition to numerous joint
walk-throughs on specific tasks ("functional" handover). [So far this
mission, Fincke has been scheduled for 7 hrs of functional handover and more
than 10 hrs of dedicated handover time. Likewise, Mike Foale has been
scheduled for over 12 hrs of functional handover and nearly 5 hrs of
dedicated handover time.]

VC6 FE-1 André Kuipers' schedule on his FD6 began with the termination,
closeout and data download (on the EGE2 laptop) of the first session of the
CIRCA experiment with its daily monitoring of blood pressure (BP) and pulse
rate in micro-G. Later, André worked his way methodically through the
"Delta" experiments ETD, ARGES, Seeds in Space, MOP, MUSCLE, and MEDIET.
[Blood pressure data transfers from the CIRCA/BMI to the EGE2 failed, and
engineers suspect a connector cable as cause. Troubleshooting is underway.
Also, the CIRCA payload annunciated an alert during crew sleep last night,
i.e., the payload was not receiving sufficient power from the U.S. "Makita"
battery. The crew tried four other "Makitas" without success. The
experiment was then transitioned to ISS power in the SM for the remainder of
the session, and it is now terminated.]

The VC6 experiment SUIT in the afternoon was deferred, to allow more time
for ARGES operations.

During today's experiment with the ARISS ham radio equipment in the SM,
Kuipers chatted with the 8-12 year old winners of the "Zeg het ISS" school
competition set up by ESA in the Netherlands and Belgium, which were
gathered at ESTEC in Noordwijk, Holland.

CDR/SO Michael Foale took Gennady Padalka and Mike Fincke through a
familiarization session with the MSS (Mobile Service System). [They first
conducted a standard review of the DOUG application, then operated the SSRMS
(Space Station Remote Manipulator System) during a two-hour period, with
both Padalka and Fincke getting hands-on experience with procedures. The
robotarm ops consisted of a single joint maneuver, followed by a joint OCAS
(operator commanded auto-sequence) maneuver, followed by another single
joint maneuver to move the arm to an MBS PDGF3 (mobile base system/power &
data grapple fixture #3) pre-grapple position and to perform a grapple and
release of the PDGF. The SSRMS was then backed off from the PDGF-3 and
stowed at the pre-grapple position. The planned camera familiarization step
was deleted to compensate for a late start; it will be accomplished as part
of normal stage operations with the MSS. Robotic ops were completed
essentially in the time allotted. Afterwards the VDS video system was
deconfigured and the UOP-DCP (utility outlet panel-to-display & control
panel) bypass power cable at the Lab RWS disconnected.]

FE Alexander Kaleri ingressed the 7S/Soyuz-213 to turn on its gas analyzer
(GA) to obtain a reading of the air composition in the spacecraft. The GA
was then powered off again.

CDR-9 Padalka performed the same activity in the new crew return vehicle

Sasha Kaleri conducted another 55-min. training session in the "Chibis" ODNT
suit as part of his preparations for returning into gravity. Mike Foale
provided assistance. [The below-the-waist reduced-pressure device ODNT (US:
LBNP) provides gravity-simulating stress to the body's
cardiovascular/circulatory system for reestablishing the body's orthostatic
tolerance (e.g., the Gauer-Henry reflex) after the six-month stay in

The new station CDR, Gennady Padalka, had half an hour reserved on his
schedule for general familiarization with his new abode in the sky. [ISS
familiarization & acclimatization is a standard daily rule for a new crew
for the first two weeks after starting station residency.]

Another functional handover task for today was a new startup by Sasha and
Gennady of the Molniya-SM/LSO payload, set up at SM window #3. The
procedure required activating the hardware including its associated EGE-1
laptop, performing test measurements for three night orbits and configuring
it for unattended use, starting today and ending 4/26. Copying of data onto
disk for return is scheduled for 4/27. [Objective of Molniya-SM, similar to
the French LSO experiment, is to record storm phenomena and other related
events in the Earth's equatorial regions. The experiment requires
orbital-sighting predictions using an up-to-date NORAD tracking TLE
(two-line element) provided by NASA. Objective of LSO is to study rare
optical phenomena occurring in the upper layers of Earth's atmosphere,
so-called "sprites" (i.e., puzzling glow phenomena observed above
thunderstorm clouds).]

Kaleri completed another sensor reading of the "Pille-MKS" radiation
dosimeter, with its ten sensors placed at various locations in the RS
(Russian segment, port cabin window, starboard cabin window, ASU toilet
facility, control panel, etc.). Pille dosage values will be returned to the

Using the SSC4 (station support computer #4) laptop in the Airlock (A/L),
Foale and Fincke initiated another maintenance charge/discharge cycle on two
EMU/spacesuit batteries (#2029 & #2030) in the A/L's BSA (battery stowage
assembly). [The charging takes about 24 hrs and will be followed by
discharge. The entire procedure is automated. Helmet light and PGT (pistol
grip tool) batteries are not involved. The EMU battery maintenance is
performed every 50 days, consisting of fully charging and then discharging
the batteries to prolong their useful life.]

Padalka worked on the Russian BIO-8 "Plazmida" experiment's Rekomb-K
hardware, changing the thermostat setting in the Kriogem-03M refrigerator to
+37 degC temperature (from +4 degC set on 4/22).

Sasha attended to the daily routine maintenance of the SOZh environmental
control & life support systems in the SM and prepared the daily "delta" file
for updating the IMS (inventory management system) database.

The Expedition 8 crew performed their regular daily physical exercise
program on TVIS treadmill and RED exerciser.

The crew reported that the Lab overhead starboard (OS) lamp housing assembly
(LHA) #4 has failed. [This LHA is directly over the EXPRESS 5 payload rack,
next to the Lab OS LHA #3, which the crew has previously reported failed.]

Update on Lab Window: The Lab window inter-pane volume ("Volume B") was
successfully evacuated yesterday as planned, but small leaks other than from
the original U-flexhose ports were detected. The Expedition 9 crew will
evacuate the window volume again after a protective cover for the new
flexhose U-jumper has been delivered on Progress 14P. [Before the
evacuation, pressure for the interstitial volume was 709 Torr (mmHg); after
the procedure, it was measured at 2.4 Torr. In addition to the evacuation,
the crew conducted inspections of the window with the Ultrasonic Leak
Detector (ULD), locating four areas where small increases in sound were
detected. These tiny leaks, which allow cabin air to reenter the evacuated
space, were marked with gray tape for further troubleshooting. At about 7
hrs after the window evacuation, the inter-pane pressure had increased to
~12.3 Torr, and a 1:55am EDT this morning it was at 28.5 Torr. The crew was
given a Go to resume the previously allowed window viewing ops (1.5
cumulative hrs max per day) with the additional restriction that the shutter
must remain closed for 8 hrs prior to the first pressure reading each day,
taken as part of the morning DPC (Daily Planning Conference) for the next 4

Update on 2-CMG Operation: (1) Analyses were performed at MCC-H from
yesterday evening until early this morning to assess CMG momentum impacts
resulting from today's robotics activities, in particular CMG desaturations
during the planned SSRMS motion. Final results predicted a momentum of less
than 80% (of the amount requiring desat) during SSRMS activity. In
actuality it peaked at only 57% during the robotics ops; also, no automatic
handover to the Russian Segment was required, and the U.S. MCS (motion
control system) maintained 2-CMG attitude control throughout the day. (2)
Tomorrow's crew schedule has been updated to include sizing two EMU suits
for the Increment 9 crewmembers and performing a communications check on one
of the suits. These activities are "get-aheads" for EMU EVA to replace the
CMG-2's RPCM (remote power controller module) S02B-D, which will be
performed at some point during Increment 9 Stage Operations.

7S/Soyuz TMA-3 (Expedition 8 + 1) flight plan:

Hatch Closing -- 4/29, 1:34pm EDT (Moscow: 4/29, 9:34pm; Astana: 4/30,
Undocking -- 4/29, 4:52pm EDT (Moscow: 4/30, 12:52am; Astana: 4/30, 3:52am)
Deorbit Burn -- 4/29, 7:21pm EDT (Moscow: 4/30, 3:21am; Astana: 4/30,
Landing -- 4/29, 8:12pm EDT (Moscow: 4/30, 4:12am; Astana: 4/30, 7:12am.

Weekly Science Update (Expedition Eight -- 23rd):

GASMAP: To the Increment 9 Crew: "Your first GASMAP activity is just around
the corner. We look forward to working with you."

Human Research Facility/Workstation (HRF WS): Continuing.

Advanced Ultrasound (ADUM): No more activities scheduled for this Increment.

Hand Posture Analyzer (HPA: Nothing new.

In-Space Soldering Investigation (ISSE): Nothing new.

Pore Formation and Mobility Investigation (PFMI): Nothing new. Space
Acceleration Measurement System (SAMS): SAMS was able to capture the Soyuz
8S docking event. The SAMS team sent a warm welcome to the Inc. 9 crew and
a fond farewell to the Inc. 8 crew. Microgravity Acceleration Measurement
System (MAMS): MAMS captured the recent docking and undocking activities.
The ground team enjoyed working with the Inc. 8 crew and looks forward to
working with Expedition 9.

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):

Protein Crystal Growth Monitoring by Digital Holographic Microscope
(PromISS): Nothing new.

Investigating the Structure of Paramagnetic Aggregates from Colloidal
Emulsions (InSPACE): Planned.

Binary Colloidal Alloy Test-3 (BCAT-3): Nothing new.

Renal Stone (RS): Thanks were uplinked to Foale and Kaleri for their
successful work with the Renal Stone experiment this past weekend. The data
they provided via IMS was "excellent". Also, the pictures taken on 4/19
were "just what the Experiment team was looking for".

Synchronized Position Hold, Engage, Reorient, Experimental Satellites
(SHERES): No more activities scheduled for this Increment.

Foot/Ground Reaction Forces during Space Flight (FOOT): Nothing new.

Materials ISS Experiment (MISSE): In progress. Deployed outside on the U.S.
Airlock. Nominal and collecting data.

Cellular Biotechnology Support Systems-Fluid Dynamics Investigation
(CBOSS-FDI): Nothing new.

Yeast Group Activation Packs (Yeast GAP): Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM): Nothing new.

Earth Science Toward Exploration Research (ESTER): No more activities
scheduled for this increment.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2): Planned.

Commercial Generic Bioprocessing Apparatus (CGBA): Nothing new.

Miscible Fluids in Microgravity (MFMG): Four runs are remaining; they will
be performed on Increment 9.

Educational Payload Operations (EPO): The educational demonstrations
performed on this increment are being edited for release and distribution to
schools around the country.

Crew Earth Observations (CEO): While no imagery was selected this week for
publication, the early CEO iceberg imagery, particularly the one with
meltponding on the surface was of great interest to the NSIDC (National Snow
and Ice Data Center). The Principal Investigator was particularly excited
that this phenomenon was noticed and documented by the crew of the ISS, and
he hopes to be able to thank them personally. Observers have noted a number
of fine city views in recent CEO imagery and a very nice sequence of the
Ganges River Delta. This is the final weekly status report to the
Expedition 8 crew. They have acquired a large amount of imagery for
investigators, who are looking forward to the opportunity to review it with
the crew in the near future.


Jacques :-)



Thread Tools
Display Modes

Posting Rules
You may not post new threads
You may not post replies
You may not post attachments
You may not edit your posts

vB code is On
Smilies are On
[IMG] code is On
HTML code is Off
Forum Jump

Similar Threads
Thread Thread Starter Forum Replies Last Post
ISS Status Report No. 17 - 2004 Jacques van Oene Space Station 0 April 4th 04 10:11 AM
ISS Status Report No. 10 - 2004 Jacques van Oene Space Station 1 February 24th 04 02:31 AM
ISS Status Report No. 06 - 2004 Jacques van Oene Space Station 0 January 29th 04 05:50 PM
ISS Status Report No. 01 - 2004 Jacques van Oene Space Station 0 January 4th 04 02:43 PM
Ed Lu Letter from Space #6 Jacques van Oene Space Station 0 July 4th 03 11:10 AM

All times are GMT +1. The time now is 09:07 PM.

Powered by vBulletin® Version 3.6.4
Copyright ©2000 - 2023, Jelsoft Enterprises Ltd.
Copyright ©2004-2023 SpaceBanter.com.
The comments are property of their posters.